# Localization phenomena in the random XXZ spin chain

### v1

@@ Abstract.
It is shown that the infinite random Heisenberg XXZ spin-$\frac{1}{2}$ chain exhibits, with probability one, spectral, eigenstate, and weak dynamical localization in an arbitrary (but fixed) energy interval in a non-trivial parameters range. The crucial step in the argument is a proof that if the Green functions for the associated finite systems Hamiltonians exhibit certain (volume-dependent) decay properties in a fixed energy interval, then the infinite volume Green function decays in the same interval as well. The pertinent finite systems decay properties for the random XXZ model had been previously verified by the authors.
@@ 1.1. The model and (informal) main result
A quantum model is associated with a Hamiltonian (self-adjoint operator) on the state space (Hilbert space). Primary goals include studying its spectrum, eigenstates, and associated dynamics.
@@
Spin chains are arrays of spins indexed by subsets $\Lambda\subset\mathbb{Z}$. If $\Lambda$ is finite, the corresponding state space is the tensor product Hilbert space $\mathcal{H}_{\Lambda}=\otimes_{i\in\Lambda}\mathcal{H}_{i}$, where each $\mathcal{H}_{i}$ is a copy of $\mathbb{C}^{2}$. For infinite $\Lambda$, we let $\mathcal{H}_{\Lambda,0}$ be the vector subspace of $\bigotimes_{i\in\Lambda}\mathcal{H}_{i}$ spanned by tensor products of the form $\bigotimes_{i\in\mathbb{Z}}\varphi_{i}$, $\varphi_{i}\in\left\{\uparrow\rangle_{i},\downarrow\rangle_{i}\right\}$, with a finite number of spin-downs, equipped with the tensor product inner product, and let $\mathcal{H}_{\Lambda}$ be its Hilbert space completion. A single spin operator $\sigma^{\sharp}$ acting on the $i$-th spin is lifted to $\mathcal{H}_{\Lambda}$ by identifying it with $\sigma^{\sharp}\otimes\mathds{1}_{\Lambda\setminus\left\{i\right\}}$, where $\mathds{1}_{\Lambda\setminus\left\{i\right\}}$ denotes the identity operator on $\mathcal{H}_{\Lambda\setminus\left\{i\right\}}$, i.e., it acts non-trivially only in the tensor product’s $i$-th component. To stress the $i$-th dependence, we will denote such single spin operator by $\sigma^{\sharp}_{i}$. More generally, If $S\subset\Lambda$, and $A_{S}$ is an operator on $\mathcal{H}_{S}$, we often identify it with its natural embedding on $\mathcal{H}_{\Lambda}$, namely with $A_{S}\otimes\mathds{1}_{\Lambda\setminus S}$.
The original motivation to study quantum spin systems goes back to the 1920s when their usefulness in explaining ferromagnetism was realized by Lenz, Ising, Dirac, and Heisenberg, among others. They are now playing a role in explaining various phenomena in a plethora of physics, computer science, chemistry, and biology topics. The rich structure associated with these systems is related to their complexity: While a single spin has a very simple state space, the dimensionality of $n$ spins grows exponentially fast with $n$. On the flip side, even for modestly sized spin systems where $n$ ranges in dozens, the computational cost of their numerical analysis is prohibitive. This problem is colloquially known in physics as the curse of dimensionality and is the main cause of our very limited theoretical understanding of such models, especially of their thermodynamic limit $\Lambda\to\mathbb{Z}$.
In this work, we study spectral and dynamical properties of the random XXZ quantum spin-$\frac{1}{2}$ chain. The random Hamiltonian $H^{\Lambda}=H_{\omega}^{\Lambda}$ on $\mathcal{H}_{\Lambda}$ is given by111 Our definition of $H^{\Lambda}$ incorporates a choice of boundary condition if $\Lambda\neq\mathbb{Z}$
@@
where $\sigma^{\pm}=\frac{1}{2}(\sigma^{x}\pm i\sigma^{y})$ are called the ladder operators and $\mathcal{N}=\tfrac{1}{2}(\mathds{1}_{\mathbb{C}^{2}}-\sigma^{z})$ is known as the number operator. The constant $\Delta$ is the anisotropy parameter; we assume $\Delta>1$ (the Ising phase). The parameter $\lambda>0$ determines the strength of a random transversal field $\sum_{i\in\Lambda}\omega_{i}\mathcal{N}_{i}$, where $\omega=\left\{\omega_{i}\right\}_{i\in\mathbb{Z}}$ is a family of random variables. Throughout this work we assume that $\left\{\omega_{i}\right\}_{i\in\mathbb{Z}}$ are independent, identically distributed random variables, whose common probability distribution $\mu$ is absolutely continuous with a bounded density and satisfies $\left\{0,1\right\}\subset\operatorname{supp}\mu\subset[0,1]$.
@@
$H^{\Lambda}$ is a well defined positive bounded self-adjoint operator for finite sets $\Lambda$. For infinite sets $H^{\Lambda}$ is understood as an unbounded positive self-adjoint operator on $\mathcal{H}_{\Lambda}$. Alternatively,, one can exploit the fact that $H^{\Lambda}$ commutes with the total number of particles operator $\mathcal{N}^{\Lambda}=\sum_{i\in\Lambda}\mathcal{N}_{i}$ to represent it as a direct sum of bounded Hamiltonians of systems with a fixed number $N$ of particles. The corresponding $N$-particles Hamiltonian $H^{\Lambda}_{N}$ is a random Schrödinger operator on a certain subgraph of $\Lambda^{N}$.
@@
The free ($\lambda=0$) XXZ system is a special variant of the famous Heisenberg model. For $\Lambda=\mathbb{Z}$, its spectrum can be determined using the method, introduced by Bethe in 1931, known as the Bethe ansatz. Its ground state energy 00 is separated by a gap of size $1-\frac{1}{\Delta}$ from the rest of the spectrum, which is expected to be absolutely continuous due to the translation invariance of the underlying Hamiltonian. This feature has been verified for some energy intervals [NaSpSt, FiSt].
@@
Starting from the first decade of the new millennium, the randomized version of this operator ($\lambda>0$) has been proposed as a prototypical model for study of the many-body localization (MBL) phenomenon in solid state physics. The initial investigations (of numerical and heuristic nature) in the physics community seemed to indicate that for large values of $\lambda$ a completely different system’s behavior emerges: The spectrum becomes pure point almost surely and the system’s dynamics changes drastically, with thermalization not occurring even in the asymptotic limit of infinite system size and evolution time. Such behavior is called the MBL phase. While localization phenomenon for single-particle systems is well understood, and indeed persists in infinite volume systems for all times, it is an open question in physics whether the MBL phase does occur. It is also not clear what is its precise characterizations, see [sierant2024] for the recent review (other physics reviews on this topic include [NandHuse, AL, abanin2019]).
@@
For single-particle systems, one usually distinguishes three types of localization: Spectral, eigenstate, and dynamical localization. We will now briefly describe these forms of localization to put our results for the random XXZ model in that context; a more detailed description can be found in Appendix A. We will use the same nomenclature for the many-body case as well, but we warn the reader that single-particle localization and MBL describe different phenomena.
@@
Spectral localization for a random operator $H_{\omega}$ in a prescribed energy interval $I$ manifests itself as pure point spectrum in $I$, almost surely. This type of localization is informative only for infinite systems, but it is necessary for formulation of the subsequent types of localization in such models.
@@
Dynamical localization is the non-spreading of initially localized wave packets (in the Schrödinger picture) or of local observables (in the Heisenberg picture), in the course of the time evolution generated by $H_{\omega}$. Eigenstate localization is typically not sufficient to guarantee dynamical localization. The decay of the eigencorrelator can be seen as a weak (in the operator-topological sense) form of dynamical localization. Since local observables for a single-particle system are either compact (in the discrete case) or relatively compact (in the continuum case), weak dynamical localization implies dynamical localization for such systems. As the result, for single-particle models the decay of the eigencorrelator is essentially synonymous with dynamical localization. This is no longer the case for many-body systems, where local observables are full rank operators.
The existing mathematical results for few-particles systems (e.g., [CS, AW2, KlN1, KlN2]) show that for sufficiently large parameters $\Delta$ and $\lambda$ the infinite volume Hamiltonian $H_{N}$ obtained by restricting the full Hamiltonian to the $N$ particle sector is spectrally, eigenstate, and weakly dynamically localized for all energies, provided $N\leq N_{0}(\Delta,\lambda)<\infty$. These methods could not be significantly improved by considering energies in a fixed interval $[0,E_{0}]$.
Our main result, stated informally below, shows that the infinite volume random XXZ model is spectrally, eigenstate, and weakly dynamically localized in a fixed energy interval $[0,E_{0}]$, uniformly in $N$, as long as $\lambda\Delta^{2}$ is sufficiently large. This regime is sometimes referred in physics as zero temperature MBL. We sketch both few-particles and zero temperature localization regimes in Figure 1. The precise mathematical formulation is given in Theorem 2.1.
@@
Let $H^{\mathbb{Z}}$ be the the random XXZ Hamiltonian on $\mathcal{H}_{\mathbb{Z}}$ with parameters $\Delta>1$and $\lambda>0$. Fix the energy interval $I(E_{0})=[0,E_{0}]$, where $E_{0}>0$. Then, if $\lambda\Delta^{2}$ is sufficiently large, we have:
@@
Figure 1. A localization cartoon for the infinite volume XXZ model in strong disorder/weak interaction regimes. The blue region $A$ is the few-particles localization $N\in[0,N_{0}]$, the green region $B$ is the zero temperature localization $E\in[0,E_{0}]$ (our result). The total region of localization can be extended to include the pink sector $C$ using existing methods. The white region $D$ is currently not understood.
@@ 1.2. Relation with existing research results and open problems
As we already mentioned, despite intensive efforts in the condensed physics community in the past two decades, even the very existence of the MBL phase remains a point of debate in the physics literature. On the mathematical level, limited progress in understanding this phenomenon has been made, mostly related to the random XXZ model. As far as we are aware, Theorem 1.1 is the first result establishing localization properties for not exactly solvable infinite spin systems in this generality. That being said, from the physics perspective results of this kind constitute a clear indication of the MBL phase only if the energy intervals are allowed to grow with the system size.
@@
Up to a few years ago, rigorous MBL-related results were restricted to the class of exactly solvable models (see the review [ARNSS]). More recently, for the random XXZ model, spectral and dynamical localization in the special energy interval $[1-\frac{1}{\Delta},2\left(1-\tfrac{1}{\Delta}\right))$, called the droplet spectrum, was established in [BeW, EKS1]. These results led to the validation of other important many-body features associated with MBL in the droplet spectrum, among them the exponential clustering properties for associated eigenfunctions and non-propagation of information [EKS3], and the area law for the entanglement entropy [BeW1].
@@
The finite volumes bound obtained in [EK22] for a fixed energy interval provides a critical input for the current paper, and were used to obtain dynamical (rather than weak dynamical) localization type results for finite systems in the same regime [EK24]. However, the bounds in [EK24] depend on the system’s volume, precluding any conclusions for infinite system. If the volume dependence could be suppressed, one could use these results in establishing the stability of the logarithmic light cone against generic local perturbations [TB]. It would be interesting to see whether the estimates developed in this paper could yield significantly stronger version of the result established in [EK24], but it is an open question if this volume dependence can be completely removed using these methods.
@@
Most of the MBL attributes would be achieved if one can show the existence of a quasi-local unitary $U$ such that with large probability $U^{*}HU$ is a diagonal Hamiltonian, in the sense that it commutes with every $\sigma_{i}^{z}$ operator. The implication is that $U^{*}HU$ can be represented as a sum of weighted products $\prod_{i\in X}\sigma_{i}^{z}$ over subsets $X\in\Lambda$ which are referred to as local integrals of motions (LIOM) in the physics literature. The LIOM representation was first proposed in [SPZA] as a possible mechanism explaining MBL, and became a popular physics tool for the heuristic derivation of the majority of MBL features. In particular, the existence of such $U$ implies that one can construct an eigenbasis for $H$ consisting of vectors of the form $U\psi$, where $\psi$ is a product state.
@@
For the Anderson model in the strong disorder regime, in any finite volume one can construct a (semi-uniformly) quasi-local $U$ that diagonalizes the Hamiltonian. (This follows from the results in [EK].) The construction relies on eigenstate localization for all energies and the possibility to label the underlying eigenfunctions according to the spatial position of their localization centers. Such labeling is attainable for the Anderson model due to the fact that one can show that with large probability the spectrum of $H$ is level spaced for sufficiently regular distribution of the random potential. The methods used in the present work are not sufficient to establish LIOM localization for two reasons: (a) We can only prove localization in a fixed energy interval and not on the whole spectrum of $H$; and (b) It is not known whether the spectrum of $H$ is level spaced with large probability. It is not clear whether it is reasonable (even from the physics perspective) to expect that such $U$ exists in the first place.
@@
The recent preprint [DGHP] considers a weak deterministic perturbation of a diagonal (and thus exactly solvable) random model, namely the Ising model in a random longitudinal field, on intervals $\Lambda$ of size $\gamma^{-c}$, where $\gamma$ is the perturbation strength and $c>0$ is a small but nonzero exponent. One of the main results announced there is the construction of a uniformly quasi-local unitary $U$ such that $U^{*}HU$ is a diagonal operator. The authors exploit it to extract a useful information about this spin chain consistent with MBL, namely that the spin transport is anomalous in this system. It will be interesting to consider the analogue of this scaling type result in the XXZ setting. Namely, one would want to consider $\lambda$ fixed and $\Delta$ large, and investigate a possible existence of a quasi-local $U$ that diagonalizes $H$ on scales $|\Lambda|\sim\Delta^{c}$.
@@
From the technical point of view, the proof of Theorem 1.1 uses the finite volume results developed in [EK22] as the input to obtain results for infinite systems, suppressing the volume dependence in these estimates. In the next section we introduce the necessary technical notation, state the precise formulation of Theorem 1.1, and present the technical results that establish it.

### v3

@@ Abstract.
It is shown that the infinite random Heisenberg XXZ spin-$\frac{1}{2}$ chain exhibits localization phenomena, such as spectral, eigenstate, and weak dynamical localization, in an arbitrary (but fixed) energy interval in a non-trivial region of the parameter space. This region depends only on the energy interval and includes weak interaction and strong disorder regimes. The crucial step in the argument is a proof that if the Green functions for the associated finite systems Hamiltonians exhibit certain (volume-dependent) decay properties in a fixed energy interval, then the infinite volume Green function decays in the same interval as well. The pertinent finite systems decay properties for the random XXZ spin chain had been previously verified by the authors.
@@ 1.1. The model and (informal) main result
Spin chains are arrays of spins indexed by subsets $\Lambda\subset\mathbb{Z}$. If $\Lambda$ is finite, the corresponding state space is the tensor product Hilbert space $\mathcal{H}_{\Lambda}=\otimes_{i\in\Lambda}\mathcal{H}_{i}$, where each $\mathcal{H}_{i}$ is a copy of $\mathbb{C}^{2}$. For infinite $\Lambda$, we let $\mathcal{H}_{\Lambda,0}$ be the vector subspace of $\bigotimes_{i\in\Lambda}\mathcal{H}_{i}$ spanned by tensor products of the form $\bigotimes_{i\in\mathbb{Z}}\varphi_{i}$, $\varphi_{i}\in\left\{\uparrow\rangle_{i},\downarrow\rangle_{i}\right\}$, with a finite number of spin-downs, equipped with the tensor product space inner product, and let $\mathcal{H}_{\Lambda}$ be its Hilbert space completion. A single spin operator $\sigma^{\sharp}$ acting on the $i$-th spin is lifted to $\mathcal{H}_{\Lambda}$ by identifying it with $\sigma^{\sharp}\otimes\mathds{1}_{\Lambda\setminus\left\{i\right\}}$, where $\mathds{1}_{\Lambda\setminus\left\{i\right\}}$ denotes the identity operator on $\mathcal{H}_{\Lambda\setminus\left\{i\right\}}$, i.e., it acts non-trivially only in the tensor product’s $i$-th component. To stress the $i$-th dependence, we will denote this single spin operator by $\sigma^{\sharp}_{i}$. More generally, if $S\subset\Lambda$, and $A_{S}$ is an operator on $\mathcal{H}_{S}$, we often identify it with its natural embedding on $\mathcal{H}_{\Lambda}$, namely with $A_{S}\otimes\mathds{1}_{\Lambda\setminus S}$.
The original motivation to study quantum spin systems goes back to the 1920s when their usefulness in explaining ferromagnetism was realized by Lenz, Ising, Dirac, and Heisenberg, among others. They are now playing a role in explaining various phenomena across physics, computer science, chemistry, and biology. The rich structure associated with these systems is related to their complexity: While a single spin has a very simple state space, the dimensionality of $n$ spins grows exponentially fast with $n$. Even for modestly sized spin systems where $n$ ranges into the dozens, the computational cost of their numerical analysis is prohibitive. This problem is colloquially known in physics as the curse of dimensionality and is the main cause of our very limited theoretical understanding of such models, especially of their thermodynamic limit $\Lambda\to\mathbb{Z}$.
In this work, we study spectral and dynamical properties of the random XXZ quantum spin-$\frac{1}{2}$ chain. The random Hamiltonian $H^{\Lambda}=H_{\omega}^{\Lambda}$ on $\mathcal{H}_{\Lambda}$ is given by [^1] Our definition of $H^{\Lambda}$ incorporates a choice of boundary condition if $\Lambda\neq\mathbb{Z}$ .
@@
where $\sigma^{\pm}=\frac{1}{2}(\sigma^{x}\pm i\sigma^{y})$ are the ladder operators and $\mathcal{N}=\tfrac{1}{2}(\mathds{1}_{\mathbb{C}^{2}}-\sigma^{z})$ is known as the number operator. The constant $\Delta$ is the anisotropy parameter; we assume $\Delta>1$ (the Ising phase). The parameter $\lambda>0$ determines the strength of a random transversal field $\sum_{i\in\Lambda}\omega_{i}\mathcal{N}_{i}$, where $\omega=\left\{\omega_{i}\right\}_{i\in\mathbb{Z}}$ is a family of random variables. Throughout this work we assume that $\left\{\omega_{i}\right\}_{i\in\mathbb{Z}}$ are independent, identically distributed random variables, whose common probability distribution $\mu$ is absolutely continuous with a bounded density and satisfies $\left\{0,1\right\}\subset\operatorname{supp}\mu\subset[0,1]$.
@@
$H^{\Lambda}$ is a well defined positive bounded self-adjoint operator for finite sets $\Lambda$. For infinite sets $H^{\Lambda}$ is understood as an unbounded positive self-adjoint operator on $\mathcal{H}_{\Lambda}$. Alternatively, one can exploit the fact that $H^{\Lambda}$ commutes with the total number of particles operator $\mathcal{N}^{\Lambda}=\sum_{i\in\Lambda}\mathcal{N}_{i}$ to represent it as a direct sum of bounded Hamiltonians of systems with a fixed number $N$ of particles. The corresponding $N$-particles Hamiltonian $H^{\Lambda}_{N}$ can be seen as a random Schrödinger operator on a certain subgraph of $\Lambda^{N}$.
@@
The free ($\lambda=0$) XXZ system is a special variant of the famous Heisenberg model. For $\Lambda=\mathbb{Z}$, its spectrum can be determined using the Bethe ansatz, a method introduced by Bethe in 1931. Its ground state energy 0 is separated by a gap of size $1-\frac{1}{\Delta}$ from the rest of the spectrum, which is expected to be absolutely continuous due to the translation invariance of the underlying Hamiltonian. This feature has been verified for some energy intervals [20, 16].
@@
Starting from the first decade of the new millennium, the randomized version of this operator ($\lambda>0$) has been proposed as a prototypical model for the study of many-body localization (MBL) phenomena in solid state physics. The initial investigations (of numerical and heuristic nature) in the physics community seemed to indicate that for large values of $\lambda$ a completely different system’s behavior emerges: The spectrum becomes pure point almost surely and the system’s dynamics changes drastically, with thermalization not occurring even in the asymptotic limit of infinite system size and evolution time. This behavior is called the MBL phase. While the localization phenomenon for single-particle systems is well understood, and indeed persists in infinite volume systems for all times, it is an open question in physics whether the MBL phase does occur. It is also not clear what its precise characterization is, see [24] for the recent review (other physics reviews on this topic include [21, 5, 1]).
@@
For single-particle systems, one usually distinguishes three types of localization: spectral, eigenstate, and dynamical localization. We will now briefly describe these forms of localization to put our results for the random XXZ model in that context; a more detailed description can be found in Appendix A. We will adopt the same nomenclature for the many-body case, though we caution the reader that single-particle localization and MBL describe different phenomena.
@@
Spectral localization for a random operator $H_{\omega}$ in a prescribed energy interval $I$ manifests itself as pure point spectrum in $I$, almost surely. This type of localization is informative only for infinite systems, but it is necessary for formulation of the different types of localization in such models.
@@
Dynamical localization is the non-spreading of initially localized wave packets (in the Schrödinger picture) or of local observables (in the Heisenberg picture), in the course of the time evolution generated by $H_{\omega}$. Eigenstate localization alone is typically not sufficient to guarantee dynamical localization. The decay of the eigencorrelator can be seen as a weak (in the operator-topological sense) form of dynamical localization. Since local observables for a single-particle system are either compact (in the discrete case) or relatively compact (in the continuum case), weak dynamical localization implies dynamical localization for such systems. As a result, for single-particle models the decay of the eigencorrelator is essentially synonymous with dynamical localization. This is no longer the case for many-body systems, where local observables are full rank operators.
The existing mathematical results for few-particles systems (e.g., [8, 3, 17, 18]) show that for sufficiently large parameters $\Delta$ and $\lambda$ the infinite volume Hamiltonian $H_{N}$ obtained by restricting the full Hamiltonian to the $N$ particle sector is spectrally, eigenstate, and weakly dynamically localized for all energies, provided $N\leq N_{0}(\Delta,\lambda)<\infty$. These methods could not be significantly improved by considering energies in a fixed interval $0,E_{0}]$.
Our main result, stated informally below, shows that the infinite volume random XXZ model is spectrally, eigenstate, and weakly dynamically localized in a fixed energy interval $0,E_{0}]$, uniformly in $N$, as long as $\lambda\Delta^{2}$ is sufficiently large. This regime is sometimes referred to in physics as zero temperature MBL, or, more precisely, low density MBL. We sketch both the few-particle and zero temperature localization regimes in Figure 1. The precise mathematical formulation is given in Theorem 2.1.
@@
Let $H^{\mathbb{Z}}$ be the random XXZ Hamiltonian on $\mathcal{H}_{\mathbb{Z}}$ with parameters $\Delta>1$and $\lambda>0$. Fix the energy interval $I(E_{0})=[0,E_{0}]$, where $E_{0}>0$. Then, if $\lambda\Delta^{2}$ is sufficiently large, we have:
@@
Figure 1. A localization cartoon for the infinite volume XXZ model in strong disorder/weak interaction regimes. The blue region AA is the few-particles localization $N\in[0,N_{0}]$, the green region $B$ is the zero temperature localization $E\in[0,E_{0}]$ (our result). The total region of localization can be extended to include the pink sector $C$ using existing methods. The white region $D$ is currently not understood.
We believe these are the first results of this kind for non-exactly solvable random spin chains. Previous localization results for random spin chains were restricted to either exactly solvable models (see the review [2], or the random XXZ spin chain in the droplet spectrum, a special energy interval at the bottom of the spectrum.
The dynamical behavior of a random system can be linked to two competing effects: The natural tendency of eigenfunctions to decay exponentially in the classically forbidden regions (which are numerous in say the large disorder regime) and their tendency to delocalize in the presence of resonances. For random Schrödinger operators on $\mathbb{Z}^{d}$, the number of resonances grows polynomially in the system size (with degree equal to $d$), and exponential decay wins, almost surely, causing localization. The curse of dimensionality can change this outcome already for a single particle operator on trees, e.g., on the Bethe lattice, as the number of resonances in this situation grows exponentially fast with the system size. This makes the underlying analysis quite delicate. Another (but somewhat related) adverse effect of the curse of dimensionality is related to the scarcity of randomness in the many-body setting: The number of random variables is actually the same as for the single particle model (that is, it grows only polynomially in the system size). As randomness is used to control the resonances in order to yield localization results, the new ideas were needed to control the exponentially large number of resonances in the many-body setting. We note that the scarcity of randomness has already been an issue for few-particle localization, where the number of resonances still grows only polynomially fast in the system size. Roughly speaking, the limitation of our result to the fixed energy interval is related to the fact that only in this context we can resolve both issues related to the curse of dimensionality.
@@ 1.2. Relation with existing research results and open problems
As we already mentioned, despite intensive efforts in the condensed matter physics community in the past two decades, even the very existence of the MBL phase remains a point of debate in the physics literature. On the mathematical level, limited progress has been made in understanding this phenomenon, mostly related to the random XXZ model. As far as we are aware, Theorem 1.1 is the first result that establishes localization properties for not exactly solvable infinite spin systems in this generality. That being said, from the physics perspective results of this kind would constitute a clear indication of an MBL phase only if the energy intervals were allowed to grow with the system size.
@@
Up to a few years ago, rigorous MBL-related results were restricted to the class of exactly solvable models (see the review [2]). More recently, for the random XXZ model, spectral and dynamical localization in the special energy interval $1-\frac{1}{\Delta},2\left(1-\tfrac{1}{\Delta}\right))$, called the droplet spectrum, was established in [7, 15]. These results led to the validation of other important many-body features associated with MBL in the droplet spectrum, among them the exponential clustering properties for associated eigenfunctions and non-propagation of information [14], and the area law for the entanglement entropy [6].
@@
The finite volumes bound obtained in [12] for a fixed energy interval provides a critical input for the current paper, and were used to obtain dynamical (rather than weak dynamical) localization type results, expressed as slow propagation of information, for finite systems in the same regime [13]. However, the bounds in [13] depend on the volume of the system, which excludes any conclusions for an infinite system. If the volume dependence could be suppressed, these results could be used to establish the stability of the logarithmic light cone against generic local perturbations [25]. It would be interesting to see whether the estimates developed in this paper could yield a significantly stronger version of the result established in [13], but it is an open question whether this volume dependence can be completely removed using these methods.
@@
Most of the MBL attributes would be achieved if one could show the existence of a quasi-local unitary $U$ such that with large probability $U^{*}HU$ is a diagonal Hamiltonian, in the sense that it commutes with every $\sigma_{i}^{z}$ operator. The implication is that $U^{*}HU$ can be represented as a sum of weighted products $\prod_{i\in X}\sigma_{i}^{z}$ over subsets $X\in\Lambda$ which are referred to as local integrals of motions (LIOM) in the physics literature. The LIOM representation was first proposed in [23] as a possible mechanism explaining MBL, and became a popular physics tool for the heuristic derivation of the majority of MBL features. In particular, the existence of such $U$ implies that one can construct an eigenbasis for $H$ consisting of vectors of the form $U\psi$, where $\psi$ is a product state.
@@
For the Anderson model in the strong disorder regime, in any finite volume one can construct a (semi-uniformly) quasi-local $U$ that diagonalizes the Hamiltonian. (This follows from the results in [11].) The construction relies on eigenstate localization for all energies and the possibility to label the underlying eigenfunctions according to the spatial position of their localization centers. Such labeling is attainable for the Anderson model due to the fact that one can show that with large probability the spectrum of $H$ is level spaced for sufficiently regular distribution of the random potential. The methods used in the present work are not sufficient to establish LIOM localization for two reasons: (a) We can only prove localization in a fixed energy interval and not on the whole spectrum of $H$; and (b) It is not known whether the spectrum of $H$ is level spaced with large probability. It is not clear whether it is reasonable (even from a physics perspective) to expect that such $U$ exists in the first place.
@@
The recent preprints [9, 19] consider weak deterministic perturbations of diagonal (and thus exactly solvable) models. In [9], the focus is on the one-dimensional Ising model with a random longitudinal field. The model is considered on intervals $\Lambda$ of size $\gamma^{-c}$, where $\gamma$ is the perturbation strength and $c>0$ is a small, non-zero exponent. A key finding announced in this work is the construction of a uniformly quasi-local unitary operator $U$ such that $U^{*}HU$ is a diagonal operator. The authors use this to derive information about the spin chain, consistent with Many-Body Localization (MBL), specifically that the spin transport in this system is anomalous.
The paper [19] examines a broad class of models on finite graphs given by $H=H_{0}+\sum_{r=i}^{\infty}\epsilon^{r}V_{r}$. Here, $\epsilon$ is a small parameter, $V_{r}$ is a spatially local operator supported on sets whose diameter is smaller than $r$, and $H_{0}$ is a diagonal operator formed by a sum of Pauli-Z strings. Assuming certain non-resonant conditions on $H_{0}$, the authors demonstrate that the propagation of information remains slow up to spatial scales that are powers of $\ln(\epsilon^{-1})$. In the case of the finite volume random XXZ model, we proved in [13] that there is slow propagation of information across all spatial scales in the region of localization (the finite energy interval).
Finally, let us mention that recently there has been interest in the physics community in low-energy MBL. In particular, in [26] it is shown that every eigenstate in an (extensive) energy interval at the bottom of the spectrum is localized for a deterministic finite system of error-correcting codes, settling the long-standing conjecture in the physics literature that such eigenstate localization is possible. Theorem 1.1 can be viewed as settling this conjecture for an infinite random system, although it applies to fixed (rather than extensive) energy intervals.
@@
The proof of Theorem 1.1 uses the finite volume results developed in [12] as input to obtain results for infinite systems, suppressing the volume dependence in these estimates. In the next section we introduce the necessary technical notation, state the precise formulation of Theorem 1.1, and present the technical results that establish it.